OpenStudy (anonymous):

Find the slope and the y-intercept of the line. Write a fractional answer in lowest terms y=-9/7x - 3/2

OpenStudy (anonymous):

slope=-9/7 and y-intercept=-3/2

OpenStudy (anonymous):

y-intercept =-3/2 and slope is -9/7